OpenStudy (anonymous):

Please help anyone! I will medal! Evaluate m – n when m = 3.4 and n = –8.7. A. –12.1 B. –5.3 C. 5.3 D. 12.1

OpenStudy (anonymous):

3.4-(-8.7)

OpenStudy (jango_in_dtown):

since m-n=3.4-(-8.7)=3.4+8.7=

OpenStudy (anonymous):

your not supposed to give direct answers:/

OpenStudy (anonymous):

Ah ok I got you guys thank you!, idk who to medal though xD he still gave how to do it btw

OpenStudy (anonymous):

yes i know but the code of conduct says to not give direct answers, you can say what it is but not to give DIRECT answers. so dont say abcd say the problem and answerr

OpenStudy (anonymous):

ok